OpenStudy (brydenwright):

Use the angle addition formula to find the expression for cos (a+b). What happens if you change it to cos(a+a)

jimthompson5910 (jim_thompson5910):

look at page 2 of this trig identity sheet http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

jimthompson5910 (jim_thompson5910):

look at page 2 of the `Sum and Difference Formulas` section you'll see this identity \[\Large \cos\left(\alpha \pm \beta\right) = \cos\left(\alpha\right)\cos\left(\beta\right)\mp\sin\left(\alpha\right)\sin\left(\beta\right)\]

jimthompson5910 (jim_thompson5910):

In your case, it means \[\Large \cos\left(a +b\right) = \cos\left(a\right)\cos\left(b\right)-\sin\left(a\right)\sin\left(b\right)\]

OpenStudy (mathmale):

You pretty much have to remember that "sum formula for the cosine" as well as the "sum formula for the sine." To answer the 2nd part of your question: Take Jim's latest input and replace that "b" with "a." What does cos (a + a) come out to? Note that we can write cos (a+a+ as cos 2a.

OpenStudy (lord_box):

\[\cos(a+a) = \cos(2a) = \cos^2(a) - \sin^2(a) = 2\cos^2(a)-1 = 1 - 2\sin^2(a)\]

OpenStudy (mathmale):

Yes...there's more than one correct formula for cos 2a = cos (a+a). Thx, Sir Box.